OpenStudy (anonymous):

Find the distance between each pair of points with the given coordinates. (-4, -9) & (1,3)

OpenStudy (anonymous):

Use this formula: \[\large \color{green}{Distance = \sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2}}\]

OpenStudy (anonymous):

\(x_1 = -4\) \(x_2 = 1\) \(y_1 = -9\) \(y_2 = 3\)

OpenStudy (anonymous):

ohk thnks

OpenStudy (anonymous):

Welcome..

OpenStudy (anonymous):

is the answer 169?

OpenStudy (anonymous):

is my answer right?

OpenStudy (anonymous):

Well I think I know what you did wrong....you forgot to square root that.

OpenStudy (anonymous):

\[\sqrt{(-5)^2+(-12)^2}=\sqrt{25+144}=\sqrt{169}\]